OpenStudy (anonymous):

which is a graph of the equation y= 2x -4?

OpenStudy (anonymous):

here are the graphs.

OpenStudy (anonymous):

Question2. equastion has (2,-1) as a solution? y= 2x - 1 y= y= x + 3 y= x - 3 y= -2x + 1

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

pic1 is the correct answer